theyadoreshayyy:

An observer (O) is located 900 feet from a building (B). The observer notices a helicopter (H) flying at a 49° angle of elevation from his line of sight. How high is the helicopter flying over the building? You must show all work and calculations to receive full credit.

theyadoreshayyy:

theyadoreshayyy:

@yassspoo can yu help me

YasssPoo:

beet

Florisalreadytaken:

a right triangle youre given \( \theta \) and adjacent height (x) is the opposite leg of the triangle so, what are we going to use? SOH? CAH? TOA?

theyadoreshayyy:

toa ?

Florisalreadytaken:

yes \[ \theta=49^\circ \] \[ adjacent=900 \] \[ opposite= x \] TOA \( \Rightarrow \tan \theta= \frac{opposite (x)}{adjacent} \) \[ \tan49^\circ = \frac{x}{900} \] \[ \tan49^\circ \times 900 =x \] can you work it out? use the website i told you about